Circular No. 7537

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000fc IN ANONYMOUS GALAXY M. Turatto, G. Altavilla, S. Benetti, E. Cappellaro, and A. Pastorello,

Osservatorio di Padova; and A. Clocchiatti, Pontificia Universidad Catolica (PUC) de Chile, report the discovery of a new supernova (V = 22.4) on three images taken on Nov. 17.09 UT with the 2.2-m telescope (+ WFI) at the European Southern Observatory (ESO), La Silla, in the course of the PUC-Padova Supernova Search. SN 2000fc is located at α = 0^{h}58^{m}33^{s}.58, δ = −27^{o}46′40′′.1 (equinox 2000.0), which is 0′′.3 west of the host galaxy’s nucleus. Nothing was present at this location on images taken a year earlier (cf. SN 1999ey, IAUC 7310). Comparison of a spectrum (range 598–1025 nm, resolution 1.5 nm), obtained on Nov. 26.10 with the ESO 3.6-m telescope (+ EFOSC2), with a database of well-observed nearby supernova spectra indicates that the new object is a type-Ia event; the best match is obtained with the spectrum of SN 1996X a few days past past maximum light and redshifted to 0.42. No direct measurement of the parent galaxy redshift is possible with the spectrum obtained.

V1494 AQUILAE A. Retter, Keele University; L. Cook, Concord, CA; R. Novák, Coper-

nicus Observatory; J. M. Saxton, University of Manchester; L. T. Jensen, Hundslund, Denmark; and D. Korčaková and J. Janík, Masaryk University, write: “Analysis of 31 nights of CCD photometry of V1494 Aql during 2000 June–August suggests a periodicity of 0.13467(2) day, or about twice as long as the signal reported on IAUC 7448. The full amplitude of the variation in the R band increased from 0.03 mag in June to 0.07 mag in August. The folded light curve shows a double structure with a shallower dip at phase 0.5, with about half the amplitude of the main periodicity.”

Visual magnitude estimates: Apr. 1.846 UT, 8.9 (A. Pearce, Nedlands, W. Australia); May 1.072, 10.1 (M. Lehký, Hradec Králové, Czech Republic); June 3.00, 10.0 (K. Sárneczky, Ágasvár, Hungary); July 1.731, 10.5 (S. Yoshida, Ibaraki, Japan); 29.99, 11.0 (J. G. S. Aguiar, Joaquim Egidio, S.P, Brazil); Sept. 3.00, 11.2 (J. Carvajal, Avila, Spain); Oct. 2.522, 11.2 (Pearce); Nov. 19.704, 11.5 (Lehký).

SUPERNOVA 2000ek IN ESO 153-G27 Corrigendum. On IAUC 7518, line 4, for 22′′ west read 17′′ west

2000 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green